# Splat-MOVER: Multi-Stage, Open-Vocabulary Robotic Manipulation via Editable Gaussian Splatting

### v2

@@ Abstract
We present Splat-MOVER, a modular robotics stack for open-vocabulary robotic manipulation, which leverages the editability of Gaussian Splatting (GSplat) scene representations to enable multi-stage manipulation tasks. Splat-MOVER consists of: (i) ASK-Splat, a GSplat representation that distills latent codes for language semantics and grasp affordance into the $3$D scene. ASK-Splat enables geometric, semantic, and affordance understanding of $3$D scenes, which is critical for many robotics tasks; (ii) SEE-Splat, a real-time scene-editing module using 3D semantic masking and infilling to visualize the motions of objects that result from robot interactions in the real-world. SEE-Splat creates a “digital twin” of the evolving environment throughout the manipulation task; and (iii) Grasp-Splat, a grasp generation module that uses ASK-Splat and SEE-Splat to propose candidate grasps for open-world objects. ASK-Splat is trained in real-time from RGB images in a brief scanning phase prior to operation, while SEE-Splat and Grasp-Splat run in real-time during operation. We demonstrate the superior performance of Splat-MOVER in hardware experiments on a Kinova robot compared to two recent baselines in four single-stage, open-vocabulary manipulation tasks, as well as in four multi-stage manipulation tasks using the edited scene to reflect scene changes due to prior manipulation stages, which is not possible with the existing baselines. Code for this project and a link to the project page will be made available soon.
@@
Gaussian Splatting, Robotic Grasping, Robotic Manipulation, Scene Editing
@@ I Introduction
Robotic manipulation requires spatial and semantic understanding of the scene in which a robot is operating. In particular, a robot must be able to identify the location and geometry of objects in its environment based on their semantic attributes. However, spatial and semantic scene understanding alone is often inadequate in robotic manipulation. In many cases, it is also critical to know the best place to grasp the object to perform the required task, that is, to be able to detect and localize grasp affordances on the object [1, 2]. Furthermore, for multi-stage manipulation tasks, where the pre-conditions for the next stage are met through the success of a previous stage, it is also essential to continually update the 3D scene model to reflect the objects’ motions and rearrangements due to the robot’s previous actions, to have an accurate scene representation that is useful for subsequent task stages.
Consequently, in this work, we introduce Splat-MOVER, a modular robotics stack for Multi-Stage, Open-Vocabulary Robotic Manipulation via Editable Gaussian Splatting. Splat-MOVER consists of three modules: (i) a $3$D Affordance-and-Semantic Knowledge Gaussian Splatting scene representation (ASK-Splat), (ii) a Scene-Editing-Enabled module for Gaussian Splatting scenes (SEE-Splat), and (iii) a Grasp-Generation module (Grasp-Splat) that uses ASK-Splat and SEE-Splat to plan grasps in multi-stage manipulation tasks.
ASK-Splat encodes semantic and affordance knowledge distilled from 2D foundation models in the $3$D GSplat scene representation. For semantic knowledge, we distill vision-language features from CLIP [3], a 2D foundation model for predicting image-text similarity. We leverage the feature distillation method in [4, 5] to extract pixel-wise image features from CLIP and embed them as attributes of the Gaussians during training of the GSplat representation. Further, we distill grasp affordances from a pre-trained 2D vision-based affordance model, Vision-Robotics Bridge (VRB) [6], which predicts a set of contact points given an image. VRB is trained from videos of humans conducting manipulation tasks. We post-process the resulting output of VRB to obtain dense pixel-wise grasp affordance codes, which are embedded as attributes in the ASK-Splat Gaussians during training. ASK-Splat adapts the concepts of a distilled feature field (DFF) [7] from the NeRF literature to the Gaussian Splatting (GSplat) model representation. ASK-Splat trains online from RGB images collected in a brief training phase prior to robot operation.
Leveraging ASK-Splat, SEE-Splat enables real-time scene-editing by extracting relevant objects from the ASK-Splat scene through a semantic 3D masking process, and subsequently moving these objects in $3$D, by applying a transformation specified by a function describing the desired motion. We use SEE-Splat to reflect the movements of objects made by a robot as it carries out multi-stage manipulation tasks, so that later stages of the task can build upon the results of earlier stages (e.g., moving a saucepan to the stove, then putting a fruit in the saucepan). Moreover, SEE-Splat enables visualization of the evolution of the scene prior to the execution of the robot’s plans.
Lastly, Grasp-Splat leverages the affordance and semantic codes embedded in the ASK-Splat scene to propose candidate grasp poses that are more likely to succeed. Grasp-Splat takes an open-vocabulary task command and produces 3D segmentation masks for the relevant objects using ASK-Splat’s embedded semantic codes. It then queries the pre-trained grasp proposal model GraspNet [8] to produce a set of grasp candidates for the objects, and re-ranks these candidates based on alignment with the grasp affordance codes embedded in the ASK-Splat scene. Following the grasp generation, Grasp-Splat selects and executes the best grasp to accomplish the sub-task.
These components together make up Splat-MOVER, illustrated in Figure LABEL:fig:splat_mover. Given a natural language description of a multi-stage manipulation task, Splat-MOVER utilizes the affordance-and-semantic-aware ASK-Splat scene representation to localize and mask the objects specified by the natural language prompt in the $3$D scene, providing specific initial and target configurations for the robotic manipulator at each stage of the manipulation task. Subsequently, SEE-Splat edits the ASK-Splat scene in real-time, reflecting the current state of the real-world scene, thereby enabling its use in multi-stage manipulation task. Leveraging object-specific grasp affordance codes from ASK-Splat, Grasp-Splat generates candidate grasp configurations for each object, at each stage of the manipulation task.
@@
We showcase Splat-MOVER’s effectiveness through hardware experiments on a Kinova robot, where Splat-MOVER achieves significantly improved success rates across four single-stage open-vocabulary manipulation tasks compared to two recent baseline methods: LERF-TOGO [9] and F3RM [5]. In three single-stage manipulation tasks, Splat-MOVER improves the success rate of LERF-TOGO by a factor of at least $2.4$, while achieving almost the same success rate in the last task ($95\%$ compared to LERF-TOGO’s $100\%$). Likewise, Splat-MOVER improves the success rates of F3RM by a factor ranging from about $1.2$ to $3.3$, across the four single-stage manipulation tasks. In addition, we demonstrate the superior performance of Splat-MOVER in four multi-stage manipulation tasks, where we leverage SEE-Splat to reflect the updates in the scene resulting from prior manipulation stages, a capability absent in existing baseline approaches.
@@
    We introduce ASK-Splat, a Gaussian Splatting scene model with embeded affordance and semantic codes, enabling geometric, semantic, and affordance scene understanding.
@@
    We propose SEE-Splat, a scene-editing module that uses 3D semantic masks from ASK-Splat to enable real-time editing of Gaussian Splatting scenes, to reflect the motion of objects in the scene due to the robot’s actions.
@@
    We introduce Grasp-Splat, a grasp-generation module that proposes affordance-aligned grasp candidates for specified objects, given a natural-language description of the manipulation task.
@@ II Related Work
In this section, we review relevant work in distilling language semantics into radiance fields and prior work in open-vocabulary scene-editing and robotic manipulation.
II-A Language-Embedded NeRFs and Gaussian Splatting
Neural Radiance Fields (NeRFs) represent $3$D scenes using multi-layer perceptrons (MLPs), parameterizing the color and density of each point in the scene [10]. NeRFs provide impressive high-fidelity scene renderings, even from novel views. However, to generate high-quality images, NeRFs generally require large MLPs, leading to significant training and rendering times and GPU memory usage [11, 12]. Existing NeRF methods employ dictionary/hash encodings [13, 14], and spatial discretization schemes [15, 16] to reduce the memory usage with fewer model parameters and improve the training times while maintaining or improving the rendering quality of NeRFs. Despite these advances, NeRFs still struggle to achieve real-time rendering speeds. In addition, NeRFs tend to generate floaters (non-existent geometry in free space), and fail to sufficiently represent free space in many scenes. Gaussian Splatting addresses these limitations of NeRFs, enabling real-time rendering [17]. Gaussian Splatting represents the environment using anisotropic $3$D Gaussians, which allow for fast tile-based rasterization.
Recent advances in foundation models have facilitated grounding language or semantic features into these $3$D radiance fields by distilling features extracted from pretrained foundation models, such as CLIP, to generate semantic relevancy maps through open-vocabulary queries. LERF[18] and MaskCLIP [4] propose language-embedded NeRFs that produce a 3D semantic relevancy map based on a query. LERF embeds language features from CLIP and DINO [19] across multiple scales, while MaskCLIP modifies the model architecture of CLIP to extract denser embeddings, which are distilled into the NeRF. Recent concurrent works [20, 21, 22, 23] have explored embedding language in Gaussian Splatting representations. Our contribution aligns with this nascent research area, as we incorporate language features from CLIP into Gaussian Splatting representations.
II-B Open-Vocabulary $3D$ Scene Editing
Editable, high-fidelity scene representations are essential to the success of robotics tasks. As the robot interacts and operates within an environment, the scene representation should be updated to reflect the evolution of the scene due to interactions of the robot in the scene, especially in multi-stage tasks. Here, we present relevant prior work in $3$D scene editing. CLIP-NeRF [24] presents an image or prompt-based NeRF editing method, composed of a disentangled conditional NeRF and a CLIP-driven manipulation module. In CLIP-NeRF, prompting can affect the entire scene, degrading the quality of the edited scene. Distilled Feature Fields (DFFs) [7] represents a 3D Neural Feature Field that provides query-based scene decomposition for NeRFs. DFF assigns a semantic feature to each coordinate within the 3D space. While NeRF’s MLP architecture makes it challenging to decompose the scene, DFF decomposition capability enables NeRF editing by semantically selecting regions of the space. Similar to LERF [18], DFF makes use of pre-trained feature encoders from CLIP [3], DINO [19], or LSeg [25] that are trained on extensive image and text data.
Since NeRFs represent the $3$D scene implicitly, scene-editing generally require costly transformations of each sampled point during the ray-tracing procedure in volumetric rendering of the scene. In our work, we leverage the explicit form of the Gaussian primitives to enable real-time scene-editing. To the best of our knowledge, this work represents the first instance of dynamic 3D-scene editing tailored for multi-stage robot manipulation.
II-C Visual Affordances
Visual affordances, as defined by [1], are characteristics of objects that suggest how they might be interacted with by humans or robots. This concept has been extensively studied, with past research primarily focused on creating models that learn these affordances from images [26, 27, 28]. These models have been applied to various datasets, such as those involving tools [29], indoor [30, 31] and outdoor scenes [32]. Other studies, such as those by [33, 34], have explored learning visual affordances by identifying human-object interaction points in video data. However, prior work has not examined embedding grasp affordances within $3$D scene representation, which is the focus of our work, for generating affordance-aware grasp configurations for manipulation tasks.
II-D Open-Vocabulary Robotic Grasping and Manipulation in Radiance Fields
Traditionally, robotic manipulation methods have focused on closed-set manipulation of objects, i.e., on the manipulation of prespecified objects. However, advances in foundation models have enabled the development of open-vocabulary robotic manipulation methods, where robots manipulate objects given a natural-language query [35]. Existing open-vocabulary robotic manipulation methods rely upon semantic scene representations, which distill semantic information from vision-language foundation models, such as CLIP [3], into a NeRF representation of the scene for semantic segmentation of the scene [18, 4]. For language-guided robotic manipulation, LERF-TOGO [9] builds upon the semantic NeRF representation in LERF [18], utilizing it with a deep-learned grasp generation method GraspNet [8] for grasp generation. F3RM [36] introduces a language-guided robotic manipulation method based on human demonstrations, computing an embedding for all task-relevant demonstrations, and learns a feature field mapping each point in space to a task embedding, utilizing the NeRF-based representation in [4]. Subsequently, candidate grasps for the robot are generated through an optimization procedure.
Although these methods provide results demonstrating their effectiveness, these methods still face a number of challenges. LERF-TOGO [9] requires the specification of a grasp location on an object by a human operator or a large language model, which might not be readily available. Likewise, F3RM [36] requires human demonstrations, which might be expensive to collect. In this paper, we propose to learn the affordance of each object in the scene, obviating the need for human guidance. We distill affordance from the VRB foundation model [6], which produces an affordance map and candidate contact points from $2$D images. Further, unlike existing open-vocabulary robotic manipulation methods, our proposed method utilizes a Gaussian Splatting scene representation, enabling real-time rendering and query performance.
III Gaussian Splatting
@@
We introduce notation relevant to this paper, before presenting Gaussian Splatting. We denote a $2$D feature embedding as ${f\in\mathbb{R}^{a\times b\times c}}$, where the first-two dimensions $(a,b)$ represent the spatial dimensions, while the last dimension $c$ represents the dimension of the embedding. Similarly, we denote $1$D feature embeddings as ${f\in\mathbb{R}^{c}}$. We provide a brief review of Gaussian Splatting, which we build upon to obtain a affordance-and-semantic-aware scene representation. We direct interested readers to [17] for a more in-depth discussion of Gaussian Splatting. In Gaussian Splatting, the scene is represented using $3$D Gaussians, parameterized by a mean ${\mu\in\mathbb{R}^{3}}$ and covariance ${\Sigma\in\mathbb{R}^{3\times 3}}$, with the probability density function (pdf) given by: ${p(x)=\eta\exp^{-\frac{1}{2}(x-\mu)^{\mathsf{T}}\Sigma^{-1}(x-\mu)}.}$ In practice, the unnormalized probability density function is used, dropping the normalization constant ${\eta\in\mathbb{R}}$. Further, the covariance matrix $\Sigma$ is represented as the product of a scaling (diagonal) matrix ${S\in\mathbb{R}^{3\times 3}}$ and a rotation matrix ${R\in\mathbb{R}^{3\times 3}}$ (given by a quaternion), expressing $\Sigma$ as: ${\Sigma=RSS^{\mathsf{T}}R^{\mathsf{T}}}$. An opacity parameter ${\alpha\in\mathbb{R}}$ and spherical harmonics parameters are assigned to each Gaussian. The spherical harmonics parameters enable the Gaussians to capture the view-dependent properties of the color of each point in space. Tile-based rasterization with $\alpha$-blending is utilized for rendering the Gaussians. Specifically, the $3$D Gaussians is projected to $2D$, with the projected covariance matrix given by: ${\Sigma^{\prime}=JW\Sigma W^{\mathsf{T}}J^{\mathsf{T}},}$ where ${J\in\mathbb{R}^{3}}$ denotes the Jacobian of the projective transformation (approximated by an affine approximation) and ${W\in\mathbb{R}^{3}}$ denotes the viewing transformation. The color of each pixel $C$ in the rendered image results from the point-based $\alpha$-blending procedure:
$C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (1)
where $N$ represents the number of Gaussians overlapping the pixel, $c_{i}$ is computed from the spherical harmonics representation of the color of Gaussian $i$, and $\alpha_{i}$ in (1) denotes the opacity of Gaussian $i$ multiplied by its unnormalized pdf (overloading the notation for the opacity parameter). The parameters of each Gaussian are optimized through stochastic gradient descent, leveraging the differentiability of the tile-based rasterization procedure to compute the gradients of the loss function $\mathcal{L}$ with respect to the parameters. In Gaussian Splatting, the loss function $\mathcal{L}$ comprises of a convex combination of the $\ell_{1}$-photometric loss and the structural similarity index measure (SSIM) loss [37], with: ${\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\mathrm{SSIM}}}$.

### v3

@@ Abstract
We present Splat-MOVER, a modular robotics stack for open-vocabulary robotic manipulation, which leverages the editability of Gaussian Splatting (GSplat) scene representations to enable multi-stage manipulation tasks. Splat-MOVER consists of: (i) ASK-Splat, a GSplat representation that distills semantic and grasp affordance features into the $3$D scene. ASK-Splat enables geometric, semantic, and affordance understanding of $3$D scenes; (ii) SEE-Splat, a real-time scene-editing module using 3D semantic masking and infilling to visualize the motions of objects that result from robot interactions in the real-world. SEE-Splat creates a “digital twin” of the evolving environment throughout the manipulation task; and (iii) Grasp-Splat, a grasp generation module that uses ASK-Splat and SEE-Splat to propose affordance-aligned candidate grasps for open-world objects. ASK-Splat is trained in real-time from RGB images in a brief scanning phase prior to operation, while SEE-Splat and Grasp-Splat run in real-time during operation. We demonstrate the superior performance of Splat-MOVER in hardware experiments on a Kinova robot compared to two recent baselines in four single-stage, open-vocabulary manipulation tasks and in four multi-stage manipulation tasks, using the edited scene to reflect changes due to prior manipulation stages, which is not possible with existing baselines. The project page is available at https://splatmover.github.io, and the code for the project will be made available after review.
@@
Gaussian Splatting, Robotic Manipulation, Scene Editing
@@ I Introduction
Open-world robotic manipulation requires spatial and semantic understanding of the scene in which a robot is operating. In particular, a robot must be able to identify the location and geometry of objects in its environment based on their semantic attributes. However, spatial and semantic scene understanding alone is often inadequate in robotic manipulation. In many cases, it is critical for the robot to know the best place to grasp the object to perform the required task, that is, to be able to detect and localize grasp affordances on the object [1, 2]. Furthermore, for multi-stage manipulation tasks, where the pre-conditions for the next stage are met through the success of a previous stage, it is essential to continually update the 3D scene model to reflect the objects’ motions and rearrangements due to the robot’s previous actions, to have an accurate scene representation that is useful for subsequent stages.
Consequently, in this work, we introduce Splat-MOVER, a modular robotics stack for Multi-Stage, Open-Vocabulary Robotic Manipulation via Editable Gaussian Splatting. Splat-MOVER consists of three modules: (i) a $3$D Affordance-and-Semantic Knowledge Gaussian Splatting scene representation (ASK-Splat), (ii) a Scene-Editing-Enabled module for Gaussian Splatting scenes (SEE-Splat), and (iii) a grasp generation module (Grasp-Splat) that uses ASK-Splat and SEE-Splat to plan grasps in multi-stage manipulation tasks. These components together make up Splat-MOVER, illustrated in Figure LABEL:fig:splat_mover. In a brief pre-scanning phase, we train the ASK-Splat model from posed RGB images of the workspace, simultaneously embedding CLIP and grasp affordance features from the images into the 3D scene model. Then, given a natural language prompt of a multi-stage manipulation task, Splat-MOVER uses ASK-Splat to localize and mask the objects from the prompt in the $3$D scene, providing 3D initial and target coordinates for the robotic manipulator at each stage of the manipulation task. Subsequently, SEE-Splat edits the ASK-Splat scene, reflecting the dynamic poses of the objects throughout the manipulation task. Finally, Grasp-Splat uses the pre-trained GraspNet model [3] to generate candidate grasps, which are then re-ranked based on their affordance scores from the embedded affordance features in ASK-Splat, to obtain grasp poses for each object, at each stage of the manipulation task. Point-to-point motion planning for the robot arm is accomplished with standard off-the-shelf planning tools, e.g., MoveIt [4].
@@
We showcase Splat-MOVER’s effectiveness through hardware experiments on a Kinova robot, where Splat-MOVER achieves significantly improved success rates across four single-stage open-vocabulary manipulation tasks compared to two recent baseline methods: LERF-TOGO [5] and F3RM [6]. In three single-stage manipulation tasks, Splat-MOVER improves the success rate of LERF-TOGO by a factor of at least $2.4$, while achieving similar success rates in a fourth task ($95\%$ compared to LERF-TOGO’s $100\%$). Likewise, Splat-MOVER improves the success rates of F3RM by a factor ranging from $1.2$ to $3.3$, across the four single-stage manipulation tasks. In addition, we demonstrate the performance of Splat-MOVER in four multi-stage manipulation tasks, where we leverage SEE-Splat to reflect the updates in the scene resulting from prior manipulation stages, a capability absent in existing baseline approaches.
@@
    We introduce ASK-Splat, a Gaussian Splatting scene model with embedded affordance and semantic features, enabling geometric, semantic, and affordance scene understanding.
@@
    We propose SEE-Splat, a scene-editing module that uses 3D object masks from ASK-Splat to enable real-time editing of Gaussian Splatting scenes, reflecting the motion of objects in the scene due to the robot’s actions.
@@
    We introduce Grasp-Splat, a grasp-generation module that proposes affordance-aligned grasp candidates for specified objects, leveraging the embedded affordance features in ASK-Splat.
These three modules come together to make Splat-MOVER, which takes in a natural-language description of a multi-stage manipulation task, and produces an executable motion plan to accomplish the task.
@@ II Related Work
Open-World Robotic Manipulation Advances in foundation models [7] (trained on large-scale datasets [8, 9]) have enabled the development of open-world robotic manipulation methods, where robots manipulate objects given natural-language task instructions at runtime, without being trained on those specific objects or tasks. Existing methods fall into one of two categories: (i) (“smart policy”) end-to-end visuo-motor policies based on a pre-trained large vision-language model [10, 11, 8], or (ii) (“smart map”) methods that use a rich 3D scene representation that embeds features from a vision-language foundation model, interfacing with a traditional manipulation planning stack to drive robot motion. Our method is of the “smart map” variety. Two other recent methods also take this general approach: LERF-TOGO [5] and F3RM [6]. LERF-TOGO [5] builds upon the semantic NeRF representation in LERF [12], coupled with GraspNet [3] for grasp generation. Similarly, F3RM [6] uses the NeRF-based distilled feature field from [13] to embed task-relevant features learned from human demonstrations into the 3D scene model, which are used to generate and optimize candidate grasps. We compare against LERF-TOGO and F3RM as baselines.
Despite their effectiveness, these methods still face a number of challenges. LERF-TOGO [5] requires the specification of a grasp location on an object by a human operator or a large language model, which might not be readily available. Likewise, F3RM [6] requires human demonstrations, which might be difficult to collect. Both methods rely on NeRFs as the underlying scene representation, which can be time-consuming to train and is not easily edited to represent changes in the environment (e.g., due to manipulation actions by the robot). In this paper, we embed affordance features from a pre-trained 2D affordance model VRB [14] into the 3D scene, thereby avoiding the need for human demonstrations or human specification of grasp locations. We also build our scene on the Gaussian Splatting representation, which is faster to train and render than NeRF, and easier to edit to reflect object motions. We introduce a novel GSplat scene editing algorithm specifically designed for the tabletop, object-centric edits needed in robotic manipulation.
Language-Embedded NeRFs and Gaussian Splats The core enabling technology behind the “smart map” variety of open-world manipulation methods is the neural feature field concept, which distills information from 2D data sources into a view-consistent 3D field by back-propagating through a differentiable image renderer. Neural Radiance Field (NeRF) [15] represents one of the earliest high-fidelity instance of this concept, distilling the 2D color information from RGB images into a 3D color and density field. NeRFs have been integrated into a number of robotics tasks, spanning navigation [16, 17], SLAM [18, 19, 20], and manipulation [21, 22]. Recent works have distilled semantic features from vision-language foundation models (such as CLIP [23], DINO [24], or LSeg [25]) into NeRFs [26, 27, 12], yielding a rich visual-semantic 3D scene representation. These methods were designed to highlight 3D semantic relevance to a text query (e.g., MaskCLIP [13] and LERF [12]), or to produce 3D object masks for scene editing (e.g., CLIP-NeRF [27] and Distilled Feature Fields (DFFs) [26]).
Unfortunately, adoption of these rich representations in robotics is still hampered by two key challenges: (i) slow training and rendering times, and (ii) inability to reflect dynamic scenes. Gaussian Splatting [28], a recent photorealistic volumetric scene representation, addresses these fundamental limitations, by providing faster training and rendering speeds than NeRFs, while also offering the potential for modeling dynamic scenes through real-time scene editing, as we demonstrate in this paper. Gaussian Splatting represents the environment using $3$D Gaussians, coupled with a fast rasterization-based image renderer. Some recent works [29, 30, 31, 32, 33] have explored embedding semantic features into 3D Gaussian Splatting, similarly to our work, but none of these have been used for robotic manipulation.
One key novelty in our method is that we distill grasp affordance features into the 3D field, together with CLIP semantic features. We use the pretrained VRB model [14], which infers a per-pixel grasp affordance metric from 2D RGB images, trained from videos of humans interacting with objects. Other works learn 2D grasp affordance models from still images ([34, 35, 36]) or from human-object interaction points in video data ([37, 38]), but none have distilled these 2D models into 3D fields to aid in robotic grasping. Both the semantic and affordance features in our model are crucial to grasp success, as shown in our experimental studies.
In concurrent work, GaussianGrasper [39] leverages the deep-learned model AnyGrasp [40] to generate robotic grasps from GSplat environments. However, GaussianGrasper does not embed grasp affordances or consider scene editing for multi-stage tasks, as we do. The concurrent work ManiGaussian [41] uses a GSplat to represent the robot arm itself, and predicts optimal arm actions for grasping given this representation. The method is shown in simulation only. Finally, the workshop paper [42] considers a GSplat method for tracking objects during robot manipulation motions, but does not consider planning or distilled affordances, as is our focus here.
III Preliminaries
@@
We introduce notation relevant to this paper, before presenting Gaussian Splatting. We denote a $2$D feature embedding as ${f\in\mathbb{R}^{a\times b\times c}}$, where the first-two dimensions $(a,b)$ represent the spatial dimensions, while the last dimension $c$ represents the dimension of the embedding. Similarly, we denote $1$D feature embeddings as ${f\in\mathbb{R}^{c}}$. We provide a brief review of Gaussian Splatting, which we build upon to obtain a affordance-and-semantic-aware scene representation. We direct interested readers to [28] for a more in-depth discussion of Gaussian Splatting. In Gaussian Splatting, the scene is represented using $3$D Gaussians, parameterized by a mean ${\mu\in\mathbb{R}^{3}}$ and covariance ${\Sigma\in\mathbb{R}^{3\times 3}}$. Further, the covariance matrix $\Sigma$ is represented as the product of a scaling (diagonal) matrix ${S\in\mathbb{R}^{3\times 3}}$ and a rotation matrix ${R\in\mathbb{R}^{3\times 3}}$ (given by a quaternion), expressing $\Sigma$ as: ${\Sigma=RSS^{\mathsf{T}}R^{\mathsf{T}}}$. An opacity parameter ${\alpha\in\mathbb{R}}$ and spherical harmonics parameters are assigned to each Gaussian. The spherical harmonics parameters enable the Gaussians to capture the view-dependent properties of the color of each point in space. Tile-based rasterization with $\alpha$-blending is utilized for rendering the Gaussians. Specifically, the $3$D Gaussians is projected to $2D$, with the projected covariance matrix given by: ${\Sigma^{\prime}=JW\Sigma W^{\mathsf{T}}J^{\mathsf{T}},}$ where ${J\in\mathbb{R}^{3}}$ denotes the Jacobian of the projective transformation (approximated by an affine approximation) and ${W\in\mathbb{R}^{3}}$ denotes the viewing transformation. The color of each pixel $C$ in the rendered image results from the point-based $\alpha$-blending procedure: ${C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),}$ where $N$ represents the number of Gaussians overlapping the pixel, $c_{i}$ is computed from the spherical harmonics representation of the color of Gaussian $i$, and $\alpha_{i}$ denotes the opacity of Gaussian $i$ multiplied by its unnormalized pdf (overloading the notation for the opacity parameter). The parameters of each Gaussian are optimized through stochastic gradient descent, leveraging the differentiability of the tile-based rasterization procedure to compute the gradients of the loss function $\mathcal{L}$ with respect to the parameters. In Gaussian Splatting, the loss function $\mathcal{L}$ comprises of a convex combination of the $\ell_{1}$-photometric loss and the structural similarity index measure (SSIM) loss [43], with: ${\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\mathrm{SSIM}}}$.